OpenStudy (anonymous):

lim tan(pi/4+h)-1/(h) h->0

OpenStudy (anonymous):

What math is this

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

Ohhh ok I'm in Math Models

OpenStudy (anonymous):

\[\tan(\frac{ \pi }{ 4+h })\] or \[\tan(\frac{ \pi }{4 }+h)\]

OpenStudy (anonymous):

@cheli320